OpenStudy (anonymous):

w^2-6w=27 I thought the answer was -3,9 but I was incorrect can someone help me

OpenStudy (anonymous):

w^2 - 6w - 27 = 0 (x+3)(x-9) = 0 You're correct!

OpenStudy (anonymous):

when I put that answer in my online homework it said it was incorrect should I have written it like you did with the ( )

OpenStudy (anonymous):

What is the entire problem? If it's a word problem, it's likely that w represents some quantity that can't reasonably have a negative value, so the -3 would be considered an extraneous solution leaving only 9 as a correct answer.

OpenStudy (anonymous):

solving equations by factoring and problem solving

OpenStudy (anonymous):

Are you sure you're providing the exact equation from the problem?

OpenStudy (anonymous):

the problem is w^2-6w=27

OpenStudy (anonymous):

it says solve the equation

OpenStudy (anonymous):

Can you take a screenshot of the entire problem for verification purposes?

OpenStudy (anonymous):

idk how to do that sorry......

OpenStudy (anonymous):

You make it into a quadratic equation by bring the 27 over the other side. This is what it looks like: w^2-6w-27=0 Now you factor and solve for w. Two numbers that multiply to -27 and add up to -6 are -9 and 3. Because -9 * 3 = -27 and -9 + 3 = -6 Therefore, the factored form is: 0= (x-9)(x+3) From this we can conclude that, x = 9, -3. This is because since the quadratic equals zero, then one of the brackets will need to be 0. This can only be if x is 9 or -3. Therefore, x = 9, -3.

OpenStudy (anonymous):

thank you I will try to put it in again